A straight rod of length L is made of a material having mass per unit length m(x)=λ|x| where x is measured from the centre of rod.The moment of inertia about an axis perpendicular to the rod and passing through one end of the rod will be? L=1m and λ=16kg/m^(2)
Ans (32/3) kg/m²
1
Expert's answer
2020-11-27T07:18:49-0500

Moment of inertia of a point mass at a distance x from the axis of rotation (assume it is y-axis) is


"\text{d}I=m(x)\text{d}x^2=\lambda(|x|)\text{d}x^2."

The total moment of inertia is


"I=\int^{0.5}_{-0.5}\lambda(|x|)\text{d}x^2=\\\space\\
=16\int^{0.5}_{-0.5}|x|\text{d}x^2=8x^2\Big|^{0.5}_{-0.5}"